The sum of the ages of two friends is 60. If three times the age of one friend is 9 less than four times the age of the other, how old are they?
Answer by mananth(16946)

let one be x
other will be 60-x
3x=4(60-x)-9
3x=240-4x-9
7x=231
x=231/7
x= 33
the other will be 60-33= 27 years
33, 27